OpenStudy (anonymous):

Define a function f : R → R by the formula f(x) = 6x − 8. Prove that f is one-to-one. Let a, b is in R, and suppose that f(a) = f(b). Then 6a − 8 = a=

OpenStudy (anonymous):

see screenshot

terenzreignz (terenzreignz):

f(x) = 6x - 8 f(a) = (simply replace x with a) f(b) = (simply replace x with b) f(a) = f(b) 6a - 8 = ?

OpenStudy (anonymous):

oh got it! 6b-8 and b :))) thanks

terenzreignz (terenzreignz):

Awesome :)